The following information is provided: Sales price per unit £95 Variable cost per unit £50 Fixed costs £80,000 Expected sales 5,000 units Calculate the margin of safety, expressed as a percentage of forecast sales.
Answer #1

Break even sales = Fixed costs / (Sales price per unit - Variable cost per unit)

= 80000 / (95-50)

= 1778 units

Margin of safety = Sales - Break even sales = 5000 - 1778 = 3222

Margin of safety % = Margin of safety/Sales

= 3222/5000

= 64.44%